OpenStudy (anonymous):

Work the item as directed. Type the equation of the directrix and state the coordinates of the focus points for x = y2/12. F = (, ) Directrix is x =

OpenStudy (anonymous):

Since now one knows how to do this I need help with this...Solve this problem using two variables. One number is 2 times another number plus 3. Their sum is 21. Find the numbers.

OpenStudy (jdoe0001):

for a parabola of that kind, the "focus form" will be \(\bf (y-k)^2=4p(x-h)\) vertex is at the coordinates of (h, k) p = distance from the vertex to the focus point since what you have is \(\bf x = \cfrac{y^2}{12}\implies 12x =y^2\implies 12(x-0)=(y-0)^2\) can you see where the vertex is at?

OpenStudy (anonymous):

Lol... I JUST figured it out like 2 minutes ago. it's 0,0 1/4...right?

OpenStudy (jdoe0001):

the vertex is at (0,0), yes the focus is at \(\bf 4p = 12 \implies p = \cfrac{12}{4}\)

OpenStudy (anonymous):

So 3?

OpenStudy (jdoe0001):

so p =3 keep in mind that the parabola has a squared "y" variable, meaning is opening sideways, "p" is positive that means, the parabola is opening towards the left-hand-side so the focus point is at (3 , 0) the Directrix is EXACTLY the same distance from the vertex, "3", but in the opposite direction of the vertex, so the the line crosses at (-3, 0) thus x = -3

OpenStudy (anonymous):

So it's (0,0) and p=3?

OpenStudy (anonymous):

Or is it3,0

OpenStudy (jdoe0001):

yeap, because the parabola has a \(\bf y^2\) component, and thus it's opening sideways, P is positive, thus from left to right

OpenStudy (jdoe0001):

say we have 2 numbers, "a" and "b" " One number is 2 times another number" \(\bf a = 2b\) "plus 3" \(\bf a = 2b +3\) "Their sum is 21" \(\bf a + b = 21\) we know \(\bf a = \color{red}{2b+3}\) so we could say that \(\bf (\color{red}{2b + 3}) + b = 21\) solve for "b" to find "b" once you find "b", to get "a", well, recall that a = 2b + 3

OpenStudy (anonymous):

There answer is 15 and 6

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I did it before I saw your post.

OpenStudy (jdoe0001):

good :)

OpenStudy (jdoe0001):

15 and 6 are correct

OpenStudy (anonymous):

I'm still confused about the other answer... it's 3,0...3. A yes or no will do.

OpenStudy (jdoe0001):

hmm... the focus point is a point, an x-coordinate and a y-coordinate, in the "focus form" of the parabola \(\bf (y-k)^2=4p(x-h)\) P = distance from the vertex to either one if P = 3, that is, the distance from the vertex to the focus is 3, then the focus is at (3, 0) that is, if you move horizontally from (0, 0) over the x-axis, 3 units to the right, you'd end up at ( 3, 0) the Directrix is just a line, if you move from the vertex (0, 0), in the opposite direction, to the left 3 units, you'd end up at (-3, 0) but the Directrix is just a line, not a ordered pair point, so the line equation is x = -3

OpenStudy (anonymous):

Ummm thanks?

OpenStudy (jdoe0001):

still confused?

OpenStudy (jdoe0001):

anyhow... yw :)